The Johnson Performing Arts center has increased in size. The area of the new building is 42% larger than the area of the original building. If the area of the new building is 56,000 square feet, determine the area of the original building
Answer by checkley79(3341)

X+(X+.42X)=56,000
X+1.42X=56,000
2.42X=56,000
X=56,000/2.42
X=23,140.50 FT^2. FOR THE ORIGINAL SIZE.